OpenStudy (anonymous):

Multiply. two-thirds times 5 and one-half 2 and two-thirds 3 and two-thirds 4 and one-third 5 and one-third

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (cggurumanjunath):

(2/3)*(5+1/2)

OpenStudy (anonymous):

whats that can u just give answer pls

OpenStudy (cggurumanjunath):

u first find (5+1/2)

OpenStudy (anonymous):

ok

OpenStudy (cggurumanjunath):

then multiply (5+1/2) by (2/3)

OpenStudy (cggurumanjunath):

how much is (5+1/2) @MuddyRedneck ?

OpenStudy (anonymous):

5 and a half

OpenStudy (cggurumanjunath):

take lcm.

OpenStudy (cggurumanjunath):

(5+1/2)=(5*2+1)/2 isn't it ?

OpenStudy (anonymous):

ok

OpenStudy (cggurumanjunath):

(5+1/2)=((5*2+1)/2)

OpenStudy (cggurumanjunath):

tell me the answer ! (5+1/2)=(5*2+1)/2 = ?

OpenStudy (anonymous):

idk

OpenStudy (cggurumanjunath):

how much is (5*2+1) ?

OpenStudy (anonymous):

15

OpenStudy (cggurumanjunath):

?

OpenStudy (cggurumanjunath):

(5*2)=10

OpenStudy (cggurumanjunath):

(5*2+1)=10+1=11 isn't it ?

OpenStudy (anonymous):

ok

OpenStudy (cggurumanjunath):

then multiply (5+1/2) by (2/3)= now tell me ! (11/2)*(2/3)

OpenStudy (cggurumanjunath):

(11/2)*(2/3) is how much ?

OpenStudy (anonymous):

1sec

OpenStudy (anonymous):

just give answer pls

OpenStudy (cggurumanjunath):

(11/2)*(2/3)=(11*2)/(2*3)

OpenStudy (cggurumanjunath):

isn't it ?

OpenStudy (anonymous):

i guess idk

OpenStudy (cggurumanjunath):

(11/2)*(2/3)=22/6=11/3

OpenStudy (cggurumanjunath):

11/3=(3+2/3)

OpenStudy (cggurumanjunath):

use a calculator !

OpenStudy (cggurumanjunath):

so the final answer is 3 and two thirds.

OpenStudy (anonymous):

finallyy that would have been a lot easier if u just told me theres ur medal

OpenStudy (cggurumanjunath):

no ! i wanted u to understand the problem.

OpenStudy (anonymous):

ok

thomaster (thomaster):

sigh... next time you're in a test you won't have anyone to tell you the answer...

OpenStudy (cggurumanjunath):

:)